CH/Agen/93/05

Agenda

Charissa Group Meeting

To be held on: Thursday, 9 Sep 1993, 1.30 pm To be held at: York

Distribution: WNC,NMC,RAC,MF,BRF,GJG,SJAH,RAH,CDJ,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ, all by e-mail/server

Agenda

Finance Discussion (BRF,WDMR,WNC & DLW) 11 a.m. S021 Buffet Lunch (Goodricke College PDR) 12.15 p.m. Inspection of Detector Module during the Morning

1. Minutes of Meeting of 14 July 1993
2. Matters Arising / Actions
3. Update on NPC/SERC matters  (DLW)
4. Daily rate for the ANU,  (BRF)

Treatment of SERC applications

5. Progress at the ANU and plans for runs (BRF,GT,VP,WNC

NMC,DLW,WDMR,MF)

6. Electronics — update and summary  (RAH,WDMR)
7. Data Acquisition Update  (RAH,WDMR,VP)
8. New detector assemblies and gas system (DLW,JK,RAC)
9. New chamber and beamline  (JSL)
10. Longer term plans for the ANU
11. Plans for Strasbourg
12. Timetable for grant applications  (BRF,DLW)
13. Any other business
14. Date of next meeting

(suggest 17 November)

W.N. Catford (Secretary) 8 September 1993